OpenStudy (anonymous):

The function b(x) = 2x − 5 determines how many bags of dog food need to be purchased for an animal shelter, where x is the number of dogs at the shelter. The shelter manager uses m(b(x)) to find the amount of money to bring for the dog food purchase. The function m(x) = 3x + 1. Solve for how much money to bring when there are 10 dogs in the shelter. Can someone talk me through this function??

OpenStudy (♪chibiterasu):

Well we know x is the number of dogs, so simply we just replace x on both sides with the number of dogs there are. In this case there are 10, so the function is now: 10m = 2(10) - 5

OpenStudy (♪chibiterasu):

Now we just simplify. 10m = 2(10) - 5 10m = 20 - 5 10m = 15 m = 15/10 m = 1.5

OpenStudy (anonymous):

so could the answer be 15?

OpenStudy (♪chibiterasu):

Now: m(b(x))

OpenStudy (♪chibiterasu):

1.5(b(10)) 1.5(10b)

OpenStudy (♪chibiterasu):

Well we now need b os we have to solve b(x) = 2x − 5

OpenStudy (♪chibiterasu):

Exact same thing as m(x) = 2x − 5 so b = 1.5

OpenStudy (♪chibiterasu):

1.5^2 • 10 2.25 • 10 22.5

OpenStudy (♪chibiterasu):

22.5 is the answer.

OpenStudy (anonymous):

That isnt an answer choice

OpenStudy (♪chibiterasu):

What are the choices

OpenStudy (anonymous):

15 31 46 52

OpenStudy (♪chibiterasu):

Hmm

OpenStudy (♪chibiterasu):

I'm not sure what I did wrong, but the closest would be 15

OpenStudy (♪chibiterasu):

Try 15, but I may have done it wrong then.